OpenStudy (tiffanymak1996):

find the arc length of the curve y=x^(2/3) for 1<= x <= 27

OpenStudy (tiffanymak1996):

i know the formula of the arc length, just couldn't integrate!

OpenStudy (anonymous):

what was the formula for arc length?

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

\[s_{ab}=\int_{a}^{b} \sqrt{1+[f'(x)]^2} \text{d}x\]

OpenStudy (tiffanymak1996):

\[L = \int\limits_{b}^{a} (\sqrt{(1+[f'(x)]^{2}}) dx\]

OpenStudy (tiffanymak1996):

yeah, that.

OpenStudy (tiffanymak1996):

so, it becomes:

OpenStudy (anonymous):

\[\int_{1}^{27} \sqrt{1+\frac{4}{9} x^{-2/3}} \ \text{d}x\]

OpenStudy (tiffanymak1996):

yeah, that's what i got

OpenStudy (tiffanymak1996):

i just don't know how to integrate

OpenStudy (anonymous):

maybe trig sub or integration by parts...idk @Mimi_x3

OpenStudy (amistre64):

would y^3/2 be easier?

OpenStudy (mimi_x3):

maybe u-sub

OpenStudy (amistre64):

y = x^(2/3) y^(3/2) = x integrate with repect to y instead if its simpler

OpenStudy (tiffanymak1996):

oh...

OpenStudy (anonymous):

wow...nice idea

OpenStudy (amistre64):

\[y^{3/2} = x\]\[\frac32y^{1/2}=x'\] \[\int_{a}^{b}\sqrt{1+(x')^2}dy\]

OpenStudy (amistre64):

1<= x <= 27 1<= y <= 9 a b are your new a and b i believe

OpenStudy (tiffanymak1996):

thx

OpenStudy (amistre64):

youre welcome; but how to integrate sqrt(1+y^3) ... hmmm

OpenStudy (tiffanymak1996):

:)

OpenStudy (amistre64):

also try to do this in vector form, or as a polar equation ... might make it simpler as well

OpenStudy (tiffanymak1996):

u substitution?

OpenStudy (anonymous):

thats y or y^3

OpenStudy (anonymous):

?

OpenStudy (amistre64):

lol, i am sooo blind :)

OpenStudy (anonymous):

lol...:)

OpenStudy (amistre64):

\[\int \sqrt{1+ky}~dy\] \[\int (1+ky)^{1/2}~dy=\frac{2}{3k}(1+ky)^{3/2}\]

OpenStudy (tiffanymak1996):

got the answer! (85 sqrt(85) -13sqrt(13))/27

OpenStudy (amistre64):

yay!!!

OpenStudy (tiffanymak1996):

thx so much

OpenStudy (amistre64):

good luck :)